# A Markovian Framing of WaveFunctionCollapse for Procedurally Generating Aesthetically Complex Environments

### v1

@@ 1 Introduction
Accross all domains and desired path lengths, non-MDP methods (which incorporate WFC-style constraints into objective functions instead of leveraging them via an MDP) perform noticeably worse; evolving the MDP action sequence leads to faster and more consistent convergence. However, on the most difficult objectives, evolving the action sequence converges very inconsistently, hinting at exploration limitations of $\mu+\lambda$ evolution.
@@ 2.1.1 WaveFunctionCollapse
The algorithm operates in two main modes: the simple tiled model, which uses explicit adjacency rules, and the overlapping model, which automatically extracts patterns from example inputs period [16] In the overlapping model, WFC can be seen as a form of self-supervised learning that learns a distribution from minimal examples—often just one—and generates similar content by maintaining the learned adjacency patterns. While WFC excels at maintaining local adjacency constraints, it struggles with global optimization objectives that are critical for gameplay, such as ensuring level solvability or balanced resource distribution [10].
@@ 2.1.2 Enhancements to WaveFunctionCollapse
Researchers have developed numerous modifications to improve WFC’s capabilities beyond its original formulation. [17] provide a comprehensive analysis of WFC as a constraint satisfaction problem, establishing its theoretical foundations and positioning it within the broader context of PCG approaches.
To improve scalability to large environments, Nie et al. proposed Nested Wave Function Collapse (N-WFC), which decomposes the generation process into nested subproblems. This approach significantly reduces time complexity from exponential to polynomial while maintaining consistency across the generated content. For non-grid environments, researchers have developed graph-based extensions of WFC that enable content generation for arbitrary topologies, enabling applications to 3D worlds and non-uniform structures [10]. LNU introduce additional extensions to WFC including the path constraint, which enforces global connectivity between specified tiles—addressing one of WFC’s key limitations regarding global structure control. This path constraint is distinct from the path-length objective optimized in the present work in that it focuses on paths between designer-specified tiles as opposed to any two most distant tiles on the map, and only ensures the existence of such a path, rather than making guarantees about its length. This distinction is significant because optimizing path length requires global reasoning about the entire map structure, which traditional WFC cannot perform. While the path constraint ensures connectivity exists, our work specifically optimizes the longest shortest path between any two points, creating a more challenging optimization problem that must balance local tile placement decisions with their global impact on path topology. [^1] Future work could extend our approach to optimize path-lengths between specific designer-specified tiles with a simple modification to our path-length objective function. Other approaches incorporate designer-driven heuristics to steer the generation process toward specific aesthetic or functional outcomes [6].  [^2] We note that future work could attempt to optimize path-lengths between specific tiles with a simple modification to our path-length objective function. Other approaches incorporate designer-driven heuristics to steer the generation process toward specific aesthetic or functional outcomes by introducing an automatic rule system along with global, multi-layer, and distance constraints to provide designers with more control over level layouts [6]. Cheng et al. demonstrate improved manual controllability and structural similarity to human-authored levels.
@@ 2.3.1 PCGRL Framework
2.3.2 Applications and Extensions of PCGRL
Recent extensions to PCGRL include adversarial reinforcement learning for PCG (ARLPCG), which uses a generator-solver architecture where one agent creates levels while another attempts to solve them [24]. This approach creates a natural curriculum for both agents, resulting in increasingly challenging yet solvable content. Cobbe et al. use procedural generation to benchmark reinforcement learning agents, showing how procedurally generated environments can test and improve the generalization capabilities of RL systems.
@@ 3.1 Binary
The binary domain (Figure 2), originally introduced in PCGRL [3], tasks the generator with modifying an existing map composed of solid and empty tiles to increase the longest shortest path between any two empty tiles by at least 20 tiles, while ensuring full connectivity of the empty space. In contrast, our formulation requires generating valid maps entirely from scratch that satisfy prescribed path length constraints, rather than modifying preexisting layouts. We also impose a stricter requirement, penalizing deviations from the target by requiring the generated map to achieve a path length of exactly $P$. Formally, for a generated map with longest shortest path length $p$, the objective function awards a score of |pP||p-P|. While the binary domain in PCGRL comprises a simple “binary” tile-set of wall and empty tiles, we use a more sophisticated tileset in which adjacent tiles—e.g. path and grass tiles—can share straight or rounded edges, creating visually smooth transitions between tile-types.
@@ 3.2 Biome/Binary Hybrid
  • hybrid river/binary: $o_{rb}=|p-P|+o_{r}$
  • hybrid field/binary: $o_{fb}=|p-P|+o_{f}$.
@@ 4.2.1 State Representation
We define each state $s_{t}$ as the current configuration of the WFC grid at timestep $t$. For a target map of size $\ell\times w$, $s_{t}$ is represented by an $\ell\times w$ matrix in which each entry at $(x,y)$ is either:
  • A tile index $t_{i}$ (for $1\leq i\leq n_{t}$, where $n_{t}$ is the total number of tile types), indicating that the cell is collapsed
  • 1-1, indicating that the cell remains uncollapsed.
The MDP terminates when no 1-1 entries remain, yielding the final artifact.
@@ 4.2.2 Action Representation
At each timestep $t$, the action $a_{t}$ specifies which tile to collapse in a single uncollapsed cell. It is parameterized as an $n_{t}$-dimensional logit vector, with each entry corresponding to a possible tile. To ensure compliance with adjacency constraints, we mask out invalid tiles by setting their logits to zero. The selected action is then $\arg\max$ over the logits, ensuring that each collapse step is constraint-respecting. Since exactly one tile is collapsed per action, a complete map requires a sequence of $\ell\times w$ actions.
@@ 4.3 Evolving an Action Sequence
   $s\leftarrow s_{0}$, $O_{x}\leftarrow 0$, $v_{x}\leftarrow 0$
@@ 5 Results
As $P$ increases, the performance gap widens. At $P=60$, non-MDP methods converge in less than 20% of runs and exhibit extreme inefficiency when they do, with baseline requiring over 170 generations on average. Evolution 1D and 2D remain below 60 generations and achieve convergence in 84% and 72% of runs respectively, underscoring their superior scalability.
@@ 6.3 Toward Scalable and Interactive Generation
Our method operates at on consumer hardware, making it viable for map generation during the game development process. However, as problems get more complex, live evolution becomes time consuming. Further research can explore learning generalized policies that are capable of generating multiple artifacts from single training instance. For example, one could train a Reinforcement Learning policy to output/edit the tile-type logits in our 1D or 2D evolutionary genomes given varying target path-lengths as in [28], with reward equal to the score of a map collapsed via WFC over these logits. Alternatively, one could apply Imitation Learning to the data generated by the evolutionary processes in this paper, similar to [29], and/or use such Imitation Learning to jump-start the RL process outlined above. Future work could explore additional heuristics beyond path length optimization, such as requirements for the level to contain important items (e.g., treasures, keys, or strategic resources) or specific gameplay mechanics (e.g., cover points, chokepoints). These would provide additional functional objectives to test the generality of the WFC-MDP approach. Additionally, RL extensions could leverage the gym environment we have implemented to test alternative optimization methods on the WFC-MDP formulation.
@@ 7.1 Optimization Parameters
We used a fixed population size of 48 across all algorithms due to hardware constraints and to maintain a consistent measure of sample efficiency. This choice may not reflect the optimal configuration for each method. In particular, FI-2Pop’s dual-population architecture may be disproportionately affected by smaller population sizes.
@@ 9.1.1 Hyperparameter Definitions
(float) Standard deviation for truncated-normal sampling of the mutation count.
@@
Gene-wise mixing: each child gene is chosen from one parent with 50% probability.
@@
Single cut-point crossover: parents swap tails at one random index.
@@
(float) Fraction of offspring produced via crossover (baseline/FI-2Pop).
cross_or_mutate
(float) Same as cross_or_mutate_proportion, used in Action Sequence Evolution.

### v2

@@ 1 Introduction
Across all domains and desired path lengths, non-MDP methods (which incorporate WFC-style constraints into objective functions instead of leveraging them via an MDP) perform noticeably worse; evolving the MDP action sequence leads to faster and more consistent convergence. However, on the most difficult objectives, evolving the action sequence converges very inconsistently, hinting at exploration limitations of $\mu+\lambda$ evolution.
@@ 2.1.1 WaveFunctionCollapse
The algorithm operates in two main modes: the simple tiled model, which uses explicit adjacency rules, and the overlapping model, which automatically extracts patterns from example inputs [16]. In the both models, WFC can be seen as a form of self-supervised learning that learns a distribution from minimal examples—often just one—and generates similar content by maintaining the learned adjacency patterns. While WFC excels at maintaining local adjacency constraints, it struggles with global optimization objectives that are critical for gameplay, such as ensuring level solvability or balanced resource distribution [10].
@@ 2.1.2 Enhancements to WaveFunctionCollapse
Researchers have developed numerous modifications to improve WFC’s capabilities beyond its original formulation. Karth and Smith provide a comprehensive analysis of WFC as a constraint satisfaction problem, establishing its theoretical foundations and positioning it within the broader context of PCG approaches.
To improve scalability to large environments, Nie et al. proposed Nested Wave Function Collapse (N-WFC), which decomposes the generation process into nested subproblems. This approach significantly reduces time complexity from exponential to polynomial while maintaining consistency across the generated content. For non-grid environments, researchers have developed graph-based extensions of WFC that enable content generation for arbitrary topologies, enabling applications to 3D worlds and non-uniform structures [10]. LNU introduce additional extensions to WFC including the path constraint, which enforces global connectivity between specified tiles—addressing one of WFC’s key limitations regarding global structure control. This path constraint is distinct from the path-length objective optimized in the present work in that it focuses on paths between designer-specified tiles as opposed to any two most distant tiles on the map, and only ensures the existence of such a path, rather than making guarantees about its length. This distinction is significant because optimizing path length requires global reasoning about the entire map structure, which traditional WFC cannot perform. While the path constraint ensures connectivity exists, our work specifically optimizes the longest shortest path between any two points, creating a more challenging optimization problem that must balance local tile placement decisions with their global impact on path topology.  [^1] We note that future work could attempt to optimize path-lengths between specific tiles with a simple modification to our path-length objective function. Cheng et al. incorporates designer-driven heuristics to steer the generation process toward specific aesthetic or functional outcomes by introducing an automatic rule system along with global, multi‑layer, and distance constraints to provide designers with more control over level layouts [6]. Yet the approach hinges on enumerated, non-local constraints—tile caps and anchored cells, hard distance windows among entities, and layer-locking of assets—that must be re-engineered for each domain, limiting generality compared with an objective-based formulation that subsumes these preferences in a unified reward.
@@ 3.1 Binary
The binary domain (Figure 2), originally introduced in PCGRL [3], tasks the generator with modifying an existing map composed of solid and empty tiles to increase the longest shortest path between any two empty tiles by at least 20 tiles, while ensuring full connectivity of the empty space. In contrast, our formulation requires generating valid maps entirely from scratch that satisfy prescribed path length constraints, rather than modifying preexisting layouts. We also impose a stricter requirement, penalizing deviations from the target by requiring the generated map to achieve a path length of exactly $P$. Formally, for a generated map with longest shortest path length $p$, the objective function awards a score of |pP|-|p-P|. While the binary domain in PCGRL comprises a simple “binary” tile-set of wall and empty tiles, we use a more sophisticated tileset in which adjacent tiles—e.g. path and grass tiles—can share straight or rounded edges, creating visually smooth transitions between tile-types.
@@ 3.2 Biome/Binary Hybrid
  • hybrid river/binary: $o_{rb}=-|p-P|+o_{r}$
  • hybrid field/binary: $o_{fb}=-|p-P|+o_{f}$.
@@ 4.1 Direct Map Evolution
Algorithm 3 WFC-MDP: One Environment Step
Input: belief grid $G\in\{0,1\}^{\ell\times w\times n_{t}}$, adjacency rules $A$, agent action (tile logits) $a$
Output: updated grid $G^{\prime}$, reward $r$
$(x^{\star},y^{\star})\leftarrow\mathrm{FindLowestEntropyCell}(G)$
$m\leftarrow G[y^{\star},x^{\star},:]$ feasibility mask at selected cell
if $(x^{\star},y^{\star})=\varnothing\text{ or }\sum_{t=1}^{T}m_{t}=0$ then
  $r\leftarrow-1000$
end if
$p\leftarrow\text{softmax}(a)$
for all t $\in\{1,\dots,n_{t}\}$ do
  if $m[t]=0$ then
   $p[t]\leftarrow 0$
  end if
end for
$t^{\star}\leftarrow\arg\max p$
$G\leftarrow\mathrm{Collapse}(G,\ (x^{\star},y^{\star})\leftarrow t^{\star})$
$(G,\ \texttt{ok})\leftarrow\mathrm{PropagateConstraints}(G,\ A,\ (x^{\star},y^{\star}))$
if not ok then
  $r\leftarrow-1000$
else if $\mathrm{IsFullyCollapsed}(G)$ then
  $r\leftarrow\text{Objective}(G)$
end if
return (G,r)(G,\ r)
@@ 4.2.1 State Representation
We define each state $s_{t}$ as the current configuration of the WFC grid at timestep $t$. For a target map of size $\ell\times w$ with $n_{t}$ tile types, $s_{t}$ is represented by an $\ell\times w\times n_{t}$ binary tensor $G_{t}$, where each depth slice encodes the feasibility of a tile at a cell:
  • $G_{t}[x,y,i]=1$ indicates that tile type $i$ is currently feasible at cell $(x,y)$.
  • $G_{t}[x,y,i]=0$ indicates that tile type $i$ is infeasible at cell $(x,y)$.
A cell is collapsed if and only if it has a one-hot feasibility vector ($\sum_{i=1}^{n_{t}}G_{t}[x,y,i]=1$); otherwise it is uncollapsed ($\sum_{i=1}^{n_{t}}G_{t}[x,y,i]>1$). The MDP terminates when every cell is collapsed (equivalently, when all $\ell\times w$ cells have one-hot feasibility vectors), yielding the final artifact.
@@ 4.2.2 Action Representation
At each timestep $t$, the action $a_{t}$ specifies which tile to collapse in a single uncollapsed cell (Algorithm 3). It is parameterized as an $n_{t}$-dimensional logit vector, with each entry corresponding to a possible tile; these logits lead to corresponding per tile probability after softmax. To ensure compliance with adjacency constraints, we mask out invalid tiles by setting their probabilities to zero. The selected action is then $\arg\max$ over the probabilities, ensuring that each collapse step is constraint-respecting. Since exactly one tile is collapsed per action, a complete map requires a sequence of $\ell\times w$ actions.
@@ 4.3 Evolving an Action Sequence
   $s\leftarrow s_{0}$; $O_{x}\leftarrow 0$; $v_{x}\leftarrow 0$
@@ 5 Results
As $P$ increases, the performance gap widens. At $P=60$, non-MDP methods converge in less than 20% of runs and exhibit inefficiency when they do, with baseline requiring over 170 generations on average. Evolution 1D and 2D remain below 60 generations and achieve convergence in 84% and 72% of runs respectively, underscoring their superior scalability.
@@ 6.3 Toward Scalable and Interactive Generation
Our method operates on consumer hardware, making it viable for map generation during the game development process. However, as problems get more complex, live evolution becomes time consuming. Further research can explore learning generalized policies that are capable of generating multiple artifacts from single training instance. For example, one could train a Reinforcement Learning policy to output/edit the tile-type logits in our 1D or 2D evolutionary genomes given varying target path-lengths as in [26], with reward equal to the score of a map collapsed via WFC over these logits. Alternatively, one could apply Imitation Learning to the data generated by the evolutionary processes in this paper, similar to [27], and/or use such Imitation Learning to jump-start the RL process outlined above. Future work could explore additional heuristics beyond path length optimization, such as requirements for the level to contain important items (e.g., treasures, keys, or strategic resources) or specific gameplay mechanics (e.g., cover points, chokepoints). These would provide additional functional objectives to test the generality of the WFC-MDP approach. Additionally, RL extensions could leverage the gym environment to test alternative optimization methods on the WFC-MDP formulation.
@@ 7.1 Optimization Parameters
We used a fixed population size of 48 across all algorithms due to hardware constraints and to maintain a consistent measure of sample efficiency. This choice may not reflect the optimal configuration for each method. In particular, FI‑2Pop’s dual‑population architecture may be disproportionately affected by smaller population sizes.
@@ 9.1.1 Hyperparameter Definitions
(float) Standard deviation for truncated‐normal sampling of the mutation count.
@@
Gene‐wise mixing: each child gene is chosen from one parent with 50% probability.
@@
Single cut‐point crossover: parents swap tails at one random index.
@@
(float) Fraction of offspring produced via crossover.